OpenStudy (anonymous):

A car travels 100 meters in 5 seconds. What is the speed of the car? speed = distance over time; 1 mile = 1609 meters 8.9 miles per hour 16 miles per hour 20 miles per hour 44.7 miles per hour

OpenStudy (mathstudent55):

\(\dfrac{100~m}{5~sec} \) How do you convert seconds to hours?

OpenStudy (er.mohd.amir):

speed=100/5 m/s=20 m/s now convert to

OpenStudy (anonymous):

im sooo confused still

OpenStudy (er.mohd.amir):

1 meter=1/1609 miles

OpenStudy (er.mohd.amir):

1 sec=1/3600 hours

OpenStudy (mathstudent55):

Since 1 mile = 1609 m, we have the conversion factors: \(\dfrac{1609~m}{1~mile} = \dfrac{1~mile}{1609~m} = 1\) We will use the conversion factor that will cancel out meters and will leave miles. We still need to convert from seconds to hours.

OpenStudy (mathstudent55):

We can go through minutes. \(60 ~sec = 1 ~min\) gives us \(\dfrac{60~sec}{1~min} = \dfrac{1~min}{60~sec} = 1\) \(60 ~min = 1 ~hour\) gives us \(\dfrac{60~min}{1~hour} = \dfrac{1~hour}{60~min} = 1\)

OpenStudy (anonymous):

im still lost

OpenStudy (mathstudent55):

Now we just pick the conversion factors that will cancel the units we don't want and will leave the units we want.

OpenStudy (mathstudent55):

I'm explaining step by step. We already have all the conversion factors we need above. Now we start with the given part: 100 m in 5 sec. \(\dfrac{100 ~m}{5~sec} \)

OpenStudy (mathstudent55):

Now we apply each conversion factor. We need to eliminate meters and end up with miles. \(\dfrac{100 ~m}{5~sec} \times \dfrac{1~mile}{1609~m}\) meters cancels meters, and we now have miles.

OpenStudy (anonymous):

44.7 miles per hour ????

OpenStudy (mathstudent55):

Now we eliminate seconds and get minutes. \(\dfrac{100 ~m}{5~sec} \times \dfrac{1~mile}{1609~m} \times \dfrac{60~sec}{1~min}\)

OpenStudy (mathstudent55):

Now we eliminated seconds and ended up with minutes. We just have one more conversion to do, to eliminate minutes and end up with hours.

OpenStudy (mathstudent55):

\(\dfrac{100 ~m}{5~sec} \times \dfrac{1~mile}{1609~m} \times \dfrac{60~sec}{1~min} \times \dfrac{60~min}{1~hour} = 44.7 ~\dfrac{miles}{hour}\)